OpenStudy (anonymous):

nonlinear equation xsquared+ysquared=25 x+y=7 help please ?

OpenStudy (anonymous):

substitution

OpenStudy (anonymous):

x^2+y^2=25 x+y=7 x=7-y (7-y)^2+y^2=25

OpenStudy (anonymous):

you know how to solve quadratics?

OpenStudy (anonymous):

a little bit . not my strong point though .

OpenStudy (anonymous):

rewrite it so that one side equals 0, you have to expand (7-y)^2

OpenStudy (anonymous):

oh ok , that makes more sense . thanks